1. An average of 4 out of 8 customers in a restaurant, are asking for green tea with their meal.

A random sample of 8 customers is selected. Show your work and use binomial to find the probability that:

(i) Exactly 2 ask for green tea with their meal?
(ii) Less than 7 ask for green tea with their meal?
Answer #1

X ~ B ( n = 8 , P = 0.5 )

Part i)

Part ii)


P ( X < 7 ) = 1 - P ( X >= 7 )

P ( X < 7 ) = 1 - 0.0351

P ( X < 7 ) = 0.9649